OpenStudy (anonymous):

A treasure chest has a volume of x3 + 9x2 + 26x + 24 cm3 and the height is x + 4 cm. Find the polynomial that would represent the area of the bottom of the treasure chest? Explain your reasoning.

OpenStudy (anonymous):

volume is basically the area of the base times the height. All you need to do is factor an (x+4) out of the volume equation and divide it out.

OpenStudy (anonymous):

\[\left( x+4 \right)\left( x ^{2}+5x+6 \right)\div \left( x+4 \right)\]